The sum of the ages of Lois and Leonard is 29years. Lois is 3 years older than Leonard. How old are they?
Answer by ikleyn(52778)

.
Answer. 13 and 16 years.

Solution

x + y = 29     (1)
x - y =  3     (2)

-------------------------  Add the equations


2x = 29 + 3

2x = 32  =============>  x = 32/2 = 16  years is the age of the older person.


Then  y = 29 - 16 = 13 years is the age of the younger person.

Solved.